OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3 0 −3 3 2

OpenStudy (anonymous):

from the second equation you can write y as x - 3. Now plug this value of y in the first equation.

OpenStudy (anonymous):

What

OpenStudy (anonymous):

Eq 1: 4x + 2y = 12 Eq 2: x - y = 3 Add y - 3 on both sides for eq 2 This gives you: x - y + y - 3 = 3 + y - 3 same as: x - 3 = y Now replace y by x - 3 in first equation: 4x + 2(x - 3) = 12 Try to solve this

OpenStudy (anonymous):

answer?

OpenStudy (anonymous):

4x + 2x - 6 = 12 => 6x = 18 => x = 3

OpenStudy (anonymous):

thanks and i just farted it smells really bad

OpenStudy (anonymous):

i hope you understand how the solution came about!!

OpenStudy (anonymous):

ewwww it stinks